# The 4th Dimension for Scaling Model Size

### v1

@@ 2.2 Measurement of Reasoning Capability
Step 2: Train LLM on Synthetic Math Data and Evaluate the Accuracy.
Solving these questions basically requires the LLM to produce operations step-by-step to derive the value of the target variables. The number of steps is a measurement of the problem’s complexity. We generated a non-overlapping training set and validation set. The training set contains problems with operations no greater than 15. The validation set contains problems with operations equal to 15, which represents the highest difficulty that the model saw during training. During validation, we feed the problem statement into the LLM and let it generate the solution and final answer. The detailed experiment configurations are described in Section 3 the experiments.
3 Experiments
3.1 Experiment Setup
Knowledge Capacity Experiment Setup.
The random number dataset has two parameters $n$ and $k$ as we explained in Section 2.1. The $n$ is chosen based on the default vocabulary size of our transformer training pipeline and does not have a specific meaning. We use $n=50257$. The $k$ is chosen to be sufficiently large so that the dataset contains enough information entropy. Prior research Allen-Zhu & Li (2024) established that approximately 2 bits of information can be stored per model parameter. With our smallest model containing 5M parameters, we designed experiments to store approximately 10M bits of information (aligning with the 2 bits per parameter capacity), requiring a sequence length of approximately $k=640,000$ random numbers. We established four baseline models with increasing parameter counts (5M, 10M, 15M, and 20M). To investigate VLD effects, we implemented three distinct patterns on the smallest (5M) and largest (20M) baseline models: cycle (repetition of layer blocks), inverse cycle (inverse repetition of layer blocks), and sequence (repeated neighboring layers). For each pattern, we studied repetition factors of 1, 2, and 3, resulting in total computational depths of 1$\times$, 2$\times$, and 3$\times$ the original model depth. All models were trained to convergence on 8$\times$×A100 GPUs using the Adam optimizer with a learning rate of 0.0002. After training, we input the training sequence to the LLM and computed the sum of entropy of the model’s softmax outputs, then compared the total entropy of the dataset with this measured value. The difference, which is the $\Delta H$ we mentioned in Section 2.1, represents the amount of information successfully absorbed by the model.
Reasoning Capability Experiment Setup.
Our dataset is generated through the iGSM framework Ye et al. (2024) (a structured mathematical problem generation system), and comprises 500K mathematical problem samples with carefully controlled difficulty parameters. Each problem adheres to consistent complexity constraints: a maximum of 15 operational steps (the individual mathematical operations required for solution), up to 20 edges in the structural dependency graph (the connections between operations), a permutation level of 5 (determining how the narrative elements are arranged in the problem text), and comprehensive solution formatting. This standardization ensures reliable measurement of reasoning abilities rather than mere calculation speed or memorization. All problems underwent rigorous validation to ensure solution accuracy. To maintain balanced difficulty distribution, we filtered problems using solution template hash values (a metric quantifying solution pattern complexity), retaining only those with hash values below 17. Each sample includes the problem text, step-by-step solution pathway, and final answer, with most problems requiring multi-step logical reasoning to reach the correct conclusion. This carefully curated dataset provided an ideal testbed for examining how various VLD configurations affect model performance on complex reasoning tasks, allowing us to systematically isolate the impact of virtual layer depth modifications on language models’ mathematical reasoning capabilities. We constructed four models with identical per-layer parameter counts but varying native layer depths (4, 8, 12, and 16 layers, corresponding to 50M, 100M, 150M, and 200M total parameters). We then applied the same VLD patterns to the 4-layer and 16-layer models as in our memorization experiment, with multiplication factors of 1, 2, and 3. All models were trained to convergence on 8× A100 GPUs using the Adam optimizer with a learning rate of 0.00002, requiring approximately 10 hours for 4-layer models and 12 hours for 8-layer models. For evaluation, we tested on 500 synthetic iGSM data samples with exactly 15 operations, which represent the highest difficulty that the model saw during training. We calculated accuracy by comparing model-generated answers against ground truth labels.
Figure 3: Knowledge capacity (mean absorbed information entropy) of different VLD patterns. Panel (a) shows how information storage capacity increases with parameter count in standard models without VLD. Panels (b-d) demonstrate that across all three VLD patterns, the information capacity remains nearly constant regardless of effective depth for both 5M and 20M parameter models.
Figure 4: Reasoning capability of different VLD patterns. Panel (a) shows performance scaling with increasing layer count in standard models. The 16-layer base model has 200M parameters, and the 12-layer version has 150M parameters. Panels (b-d) compare performance across three VLD patterns: Sequence, Cycle, and Inverse Cycle, applied to both 4-layer 50M-parameter (blue) and 8-layer 100M-parameter (green) base models.
3.2 Experiment Findings
In this section, we present and interpret key findings derived from our carefully controlled experiments described previously. Specifically, we explore how classical scaling and VLD scaling differently influence knowledge capacity and reasoning capability. We organize our experimental results into primary and secondary observations, detailing each with corresponding evidence and insights.
Primary Finding: VLD scaling significantly increases the reasoning capability while forcing the knowledge capacity to stay almost constant.
As illustrated by Figure 4, applying VLD patterns dramatically boosts models’ reasoning performance. Notably, the accuracy of the 4-layer model increases from 46.3% to 62.05% with a cycle VLD factor of 2, marking a substantial improvement. Simultaneously, Figure 3 confirms that despite these enhancements in reasoning capability, the knowledge storage capacity remains largely unchanged. We also showed this result in Figure 1. This distinctive attribute of VLD scaling clearly differentiates it from classical model scaling, as it enables substantial reasoning improvements without significantly altering the knowledge capacity.
Secondary Finding 1: In classical model size scaling, the knowledge capacity is proportional to the number of parameters.
In Figure 3 (a), we observe a clear and consistent upward trend in the amount of information entropy absorbed as the number of parameters increases in standard (non-VLD) models. This trend aligns with established theoretical expectations and prior empirical findings (Allen-Zhu & Li, 2024), confirming that larger models indeed have greater knowledge storage capabilities. Specifically, each increment in parameter count corresponds to a proportional increase in the amount of entropy that the model can absorb, reinforcing the direct relationship between model size and knowledge capacity.
Secondary Finding 2: In VLD scaling, the knowledge capacity keeps almost constant with small variations.
Panels (b–d) of Figure 3 illustrate an intriguing phenomenon: despite applying different virtual logical depth (VLD) patterns—including cycle, inverse cycle, and sequence—to significantly increase computational depth, the models’ absorbed information entropy remains remarkably stable. Minor variations appear but lack statistical significance, suggesting that VLD scaling primarily influences computational processes rather than storage capacity. These results underline a key property of VLD scaling—maintaining a consistent knowledge capacity irrespective of the effective algorithmic depth achieved by parameter reuse.
Secondary Finding 3: It is not necessary to increase the number of parameters to improve the inherent reasoning capability of the model.
The results depicted in Figure 4 compellingly demonstrate that smaller models enhanced with VLD can surpass significantly larger baseline models in reasoning performance. For instance, the 50M-parameter model using the cycle pattern (factor 2, effective depth=12) achieves 62.05% accuracy, surpassing the native 150M-parameter model (12-layer) with only 61.15% accuracy. This remarkable outcome emphasizes the efficacy of VLD as a parameter-efficient alternative for enhancing reasoning capabilities, indicating that increased reasoning does not necessitate an expansion in model parameters.
Discussion 1: The results above seem to be universal and generally hold under the scope of our experiments.
Through extensive experimentation across multiple model configurations, various VLD patterns, and different computational depths, the observed outcomes consistently affirm that VLD scaling robustly enhances reasoning without significantly altering knowledge capacity. Figures 3 and 4 collectively reinforce this universality, showing that regardless of the particular experimental setup, the key properties of VLD scaling—stable knowledge capacity and enhanced reasoning—remain consistent. This generalizability underscores the robustness of our findings within the rigorously defined scope of our experiments.
Discussion 2: The scaling result is not 100% monotonic and has a few outliers.
While our general findings indicate consistent trends, there exist instances where scaling produced counterintuitive results. For example, Figure 4 (c) highlights a case in which increasing the effective depth to 32 with VLD unexpectedly results in lower reasoning accuracy compared to a depth of 16. This anomaly suggests the potential presence of performance bottlenecks when scaling exclusively in a single dimension, similar to findings from previous research (Wei et al., 2022) that effective scaling requires balancing across multiple dimensions such as width, depth, data, and compute. Future studies should investigate these non-monotonic behaviors more thoroughly to optimize VLD scaling strategies.
3.3 Limitations and Further Experiments
While the current experiments already show reasonable results with carefully controlled experiments, it does come with some limitations. First, the reasoning capability should not be only measured on math benchmarks, and should also include code benchmarks and general scientific reasoning. Expanding the diversity and complexity of these benchmarks is essential to more accurately assess the generality and real-world applicability of the observed improvements. Second, though the current experiments show that VLD scaling has great potential, questions remain regarding the practical integration of this strategy into the standard training pipelines for widely adopted models. The efficacy and efficiency of applying VLD to very large-scale models trained on extensive natural language corpora have not yet been established. For instance, the optimal configuration of VLD patterns—sequence, cycle, or inverse cycle—may differ when transitioning from synthetic datasets to more realistic datasets involving mixed tasks. In order to address these questions, we provide supplemental content in https://vldscaling.ngrok.io, which will facilitate deeper exploration, replication, and future advancements in VLD-based scaling strategies.
4 Related Work
Parameter Sharing and Layer Reuse.
Parameter reuse has garnered considerable attention in recent research. Reid et al. (2021) introduced Subformer, which reused 50% of parameters for compression purposes, while Liu et al. (2023) presented ESP (Efficient Shared Parameterization) that reused central tensor matrices. However, both introduced atypical elements to transformer architectures, complicating direct comparison with standard models. Bhojanapalli et al. (2021) developed Reuse Transformers, exploring attention mechanism reuse while sharing less than 10% of attention heads. Takase & Kiyono (2021) proposed three distinct layer-sharing patterns but conducted experiments only on small networks without scaling analysis. More recently, Hay & Wolf (2024) proposed Dynamic Layer Tying, employing reinforcement learning to identify optimal layer-sharing configurations. In parallel, several studies have explored Looped Transformers (Fan et al., 2024; Saunshi et al., 2025; Yang et al., 2023), which iteratively apply transformer layers to enhance model capabilities without proportionally increasing parameters. Our work builds upon these foundations by systematically analyzing VLD scaling dynamics and its distinct impact on reasoning capability versus knowledge capacity.
Measuring Reasoning and Knowledge Capacity.
Quantitative assessment of reasoning and knowledge remains challenging. For reasoning capabilities, Ye et al. (2024) employed synthetic GSM8K benchmarks to isolate inherent reasoning abilities from memorized solutions. Regarding knowledge capacity, Allen-Zhu & Li (2024) demonstrated that LLMs store approximately 2 bits of knowledge per parameter, providing a quantitative metric for model memory. Carlini et al. (2022) established methodologies for quantifying memorization in LLMs, offering insights into information retention mechanisms. Understanding how large language models (LLMs) store and retrieve information is central to evaluating their memory mechanisms. Liang et al. (2024) introduced the SAGE framework, which enhances LLM agents with reflective and memory-augmented abilities. Xu et al. (2025) propose A-Mem, an agentic memory framework designed to enhance the long-term reasoning and adaptability of LLM agents. Unlike prior memory systems that rely on static workflows or graph-based retrieval, A-Mem adopts principles from the Zettelkasten method to dynamically organize, interlink, and evolve memories.
5 Conclusion
In this study, we studied Virtual Logical Depth (VLD), a novel scaling dimension that enhances reasoning capabilities without increasing parameter counts. Our experiments demonstrate that while traditional scaling linearly increases knowledge capacity with parameter count, VLD uniquely maintains a nearly constant knowledge capacity but significantly boosts reasoning performance. Remarkably, smaller models with VLD can surpass larger standard models in complex reasoning tasks. Future research should explore combining VLD with classical scaling dimensions and validate these results across diverse reasoning tasks. Our open-source implementation is available online to facilitate further exploration and practical integration of VLD scaling.

### v3

@@ E.1 Random Sequence Data for Knowledge Capacity
E.3 Real-World Training Corpus Specifications
For the post-training experiments, we constructed a diverse multi-domain corpus totaling approximately 2.3 billion tokens across natural language, code, and mathematical reasoning domains. Following the methodology established in Shen et al. (2024), we carefully curate datasets across three critical reasoning domains with strategic subset selection to ensure balanced coverage.
Table 3: Real-World Training Corpus for Reasoning Capability Evaluation. The dataset contains totally around 2.3 billion tokens, which comes from three categories: natural language, code and math data.
Category Dataset Tokens (M)
Natural Language SFT Data xP3x 317
OpenAssistant 20.9
OpenHermes 396
UltraChat 251
Code SFT Data Magicoder-OSS 50.5
Magicoder-Evol 74.72
Code-290k-ShareGPT 159.99
CommitPackFT 38
Evol-Code Alpaca 74.77
Math Data Open-web-math 317
algebraic-stack 353.4
TemplateGSM 79.7
StackMathQA 225
OpenR1-Math-220k 46.7
Data Example
We show some training data examples for each categories below:
(1) Natural Language (example from OpenAssistant dataset):
Question: How can one fight back when a monospony had been created? Answer: Monopsony refers to a market structure where there is only one buyer of a good or service. In the context of labor markets, a monopsony occurs when there is only one employer in a particular geographic area or industry, giving them significant power to set wages and working conditions. Research has shown that the existence of monopsonies can lead to lower wages and reduced job opportunities for workers. For example, a study by the National Bureau of Economic Research found that in industries with high levels of concentration, workers earn lower wages and are less likely to receive benefits such as health insurance.
(2) Code (example from Magicoder-OSS dataset):
Problem: Write a Python function to find the longest palindromic substring. Solution: def longest_palindrome(s: str) -> str: if not s: return "" def expand_around_center(left, right): while left >= 0 and right < len(s) and s[left] == s[right]: left -= 1; right += 1 return right - left - 1 start, max_len = 0, 0 for i in range(len(s)): len1 = expand_around_center(i, i)     # odd-length len2 = expand_around_center(i, i + 1) # even-length current_max = max(len1, len2) if current_max > max_len: max_len = current_max start = i - (current_max - 1) // 2 return s[start:start + max_len]
(3) Math (example from Open-web-math dataset):
Text: Physical Quantity Analogous to Inductance 1. May 12, 2013tapan_ydv Hi, I understand that some physical quantities in electromagnetism are analogous to physical quantities in heat transfer. For instance, electric field is analogous to temperature gradient. I want to know which physical quantity in heat transfer is analogous to inductance ($L$)? Regards, 2. May 12, 2013tiny-tim welcome to pf! hi tapan_ydv! welcome to pf! i don’t know about a heat transfer analogy, but a hydraulics analogy is a paddle-wheel. A heavy paddle wheel placed in the current. The mass of the wheel and the size of the blades restrict the water’s ability to rapidly change its rate of flow (current) through the wheel due to the effects of inertia, but, given time, a constant flowing stream will pass mostly unimpeded through the wheel, as it turns at the same speed as the water flow … (from http://en.wikipedia.org/wiki/Hydraulic_analogy#Component_equivalents) 3. May 12, 2013technician In mechanics … inertia. 4. May 12, 2013tiny-tim how? 5. May 12, 2013technician Reluctance to change … as in a paddle wheel. Last edited: May 12, 2013
Appendix F Model Architectures and Training Details
F.1 GPT-2 Model Configurations for Knowledge Capacity Experiments
We established four GPT-2 models with increasing parameter counts while maintaining architectural consistency. The models are configured as follows:
  • 5M parameters: 4 layers, hidden size 92
  • 10M parameters: 4 layers, hidden size 128
  • 15M parameters: 4 layers, hidden size 156
  • 20M parameters: 4 layers, hidden size 184
For VLD experiments, we applied the three patterns (Sequence, Cycle, Inverse Cycle) to the 5M and 20M baseline models with repetition factors $\times\{1,2,3\}$.
Training Configuration: we show the training configurations in Table  4.
Table 4: Training Hyperparameters and Configuration for Knowledge Capacity Experiments
Parameter Value
Optimizer Adam
Learning rate $2\times 10^{-5}$
Precision FP16
Micro-batch size 24
Sequence length (training) 768
Sequence length (evaluation) 1024
Max position embeddings 1024
Tensor model parallel size 1
Pipeline model parallel size 1
Seed 1234
Tokenizer type RandomNumberTokenizer
Hardware 8 × A100 GPUs
F.2 GPT-2 Model Configurations for Reasoning Experiments
We constructed four GPT-2 based models with identical per-layer parameter counts but varying native layer depths:
  • 4 layers ( 50M parameters): Base architecture for VLD application
  • 8 layers ( 100M parameters): Base architecture for VLD application
  • 12 layers ( 150M parameters): Baseline comparison only
  • 16 layers ( 200M parameters): Baseline comparison only
VLD patterns were applied to 4-layer and 8-layer models with multiplication factors $\times\{1,2,3\}$ (and up to $\times 5$ for cycle pattern).
Training Configuration: we show the training configurations in Table  5.
Table 5: GPT-2 Training Configuration for Reasoning Experiments
Parameter Value
Model Architecture
Number of layers 4 (base model)
Hidden size 768
Number of attention heads 12
Sequence length 768
Max position embeddings 2048
Attention backend Auto
Training Hyperparameters
Micro batch size 48
Global batch size 768
Training iterations 30,000
Learning rate 0.0005
LR decay style Cosine
LR warmup fraction 0.019
Weight decay 0.001
Adam beta1 0.9
Adam beta2 0.95
Gradient clipping 1.0
Precision FP16
F.3 LLaMA-3.2-3B-Instruct Fine-tuning Setup
For real-world evaluation, we fine-tune both Base and Cycle VLD variants of LLaMA-3.2-3B-Instruct on the multi-domain corpus. Both models are initialized with identical pretrained weights to ensure fair comparison, with the Cycle VLD variant incorporating the layer repetition pattern before training commences. The training details are listed in Table  6.
Table 6: LLaMA-3.2-3B-Instruct Fine-tuning Configuration
Parameter Value
Base model LLaMA-3.2-3B-Instruct
Total parameters 3,218,828,288
Trainable parameters (LoRA) 6,078,464 (0.19%)
Fine-tuning method LoRA
LoRA target modules gate_proj, v_proj, k_proj, up_proj, o_proj, down_proj, q_proj
Optimizer AdamW
Precision BFloat16
DeepSpeed stage ZeRO Stage 3
World size 8 GPUs
Train batch size 192
Micro batch size per GPU 6
Gradient accumulation steps 4
Gradient clipping 1.0
Distributed training DeepSpeed ZeRO-3 + LoRA
Parameter offloading Enabled
Gradient checkpointing Enabled
Memory optimization ZeRO-3 with parameter persistence
Appendix G Evaluation Protocols and Metrics
G.1 Knowledge Capacity Measurement Methodology
After training, we input the training sequence to the LLM and compute the sum of entropy of the model’s softmax outputs, then compare the total entropy of the dataset with this measured value. The difference, $\Delta H$, represents the amount of information successfully absorbed by the model.
The mean absorbed information entropy is defined as the average amount of information absorbed per token, calculated as the reduction from the original uncertainty (15.6 bits for $n=50257$) to what the model predicts.
G.2 Reasoning Capability Assessment Protocols
In-Distribution Evaluation:
We test models on 500 synthetic iGSM samples with exactly 15 operations, matching the maximum complexity seen during training. Each problem requires multi-step mathematical reasoning, and we calculate exact-answer accuracy by comparing model-generated solutions against ground truth labels using string matching on the final numerical answers.
Out-of-Distribution Evaluation:
We assess model performance on higher complexity problems containing 20 and 21 operations, respectively. These problems maintain the same mathematical reasoning structure but require additional computational steps. All evaluations use identical decoding parameters (greedy decoding) and prompting strategies across model variants to ensure fair comparison.
G.3 Real-world Benchmark Evaluation Details
We conduct comprehensive evaluations on diverse real-world reasoning benchmarks following established evaluation protocols (Habib et al., 2023):
For evaluation metrics, we maintain consistency with prior work (Habib et al., 2023; Hendrycks et al., 2021; Rein et al., 2024), employing exact match accuracy for mathematical and scientific reasoning tasks (Math500, AIME, GPQA) and pass@1 accuracy for code generation benchmarks (HumanEval, MBPP) to measure functional correctness.
Test data examples
We show some test data examples for each categories below:
HumanEval:
Problem:
from typing import List def has_close_elements(numbers: List[float], threshold: float) -> bool: """ Check if in given list of numbers, any two numbers are closer than the given threshold. >>> has_close_elements([1.0, 2.0, 3.0], 0.5) False >>> has_close_elements([1.0, 2.8, 3.0, 4.0, 5.0, 2.0], 0.3) True """ Solution:
def has_close_elements(numbers: List[float], threshold: float) -> bool: for idx, elem in enumerate(numbers): for idx2, elem2 in enumerate(numbers): if idx != idx2: distance = abs(elem - elem2) if distance < threshold: return True return False
MBPP:
Problem: Write a function to find the longest chain which can be formed from the given set of pairs. Solution:
class Pair(object): def __init__(self, a, b): self.a = a self.b = b def max_chain_length(arr, n): best = 0 mcl = [1 for _ in range(n)] for i in range(1, n): for j in range(0, i): if arr[i].a > arr[j].b and mcl[i] < mcl[j] + 1: mcl[i] = mcl[j] + 1 for i in range(n): if best < mcl[i]: best = mcl[i] return best
Math500:
Problem: If $f(x)=\frac{3x-2}{x-2}$, what is the value of f(2)+f(1)+f(0)f(-2)+f(-1)+f(0)? Express your answer as a common fraction. Solution: $f(-2)+f(-1)+f(0)=\frac{3(-2)-2}{-2-2}+\frac{3(-1)-2}{-1-2}+\frac{3(0)-2}{0-2}=\frac{-8}{-4}+\frac{-5}{-3}+\frac{-2}{-2}=2+\frac{5}{3}+1=\boxed{\frac{14}{3}}$
AIME:
Problem: Let $x,y$ and $z$ be positive real numbers that satisfy the following system of equations: $\log_{2}\left({x\over yz}\right)={1\over 2}$ $\log_{2}\left({y\over xz}\right)={1\over 3}$ $\log_{2}\left({z\over xy}\right)={1\over 4}$ Then the value of $\left|\log_{2}(x^{4}y^{3}z^{2})\right|$ is $\tfrac{m}{n}$ where $m$ and $n$ are relatively prime positive integers. Find $m+n$. Solution: Denote $\log_{2}(x)=a$, $\log_{2}(y)=b$, and $\log_{2}(z)=c$. Then, we have: $a-b-c=\frac{1}{2}$, $a+b-c=\frac{1}{3}$, $a-b+c=\frac{1}{4}$. Now, we can solve to get $a=\frac{-7}{24},b=\frac{-9}{24},c=\frac{-5}{12}$. Plugging these values in, we obtain $|4a+3b+2c|=\frac{25}{8}\implies\boxed{033}$.
GPQA:
Problem: Two quantum states with energies $E_{1}$ and $E_{2}$ have lifetimes of $10^{-9}\,\mathrm{s}$ and $10^{-8}\,\mathrm{s}$, respectively. We want to clearly distinguish these two energy levels. Which of the following options could be their energy difference so that they can be clearly resolved? Solution: $10^{-4}\,\mathrm{eV}$
Appendix H Additional Experimental Results
H.1 Layer Selection Strategy Analysis
We also conducted ablation studies to assess how the placement of shared layers affects performance. Using an 8-layer model with cycle pattern, we compared different layer selection strategies: applying VLD to all layers (1-8), early layers only (1-4), and later layers only (5-8).
Table 7: GSM Performance of the 8-Layer Model with Varying Layer Selection Strategies
VLD Pattern No VLD 1-8 layers 1-4 layers 5-8 layers
cycle_8 60.5 62.0 54.2 64.2
Appendix I Limitations and Future Work
While our comprehensive experiments demonstrate VLD’s effectiveness across multiple configurations, several limitations constrain the scope and generalizability of our findings. The non-monotonic scaling behavior observed in our experiments, where performance occasionally decreases at higher VLD factors, warrants deeper investigation to understand the underlying mechanisms and identify potential fixed points where further scaling yields diminishing returns.
The non-monotonic behavior presents intriguing research opportunities: systematically exploring whether VLD has inherent fixed points beyond which performance no longer improves, investigating optimal repetition patterns through combination with other techniques like mixture-of-experts or sparse attention mechanisms, and developing principled methods to identify the best VLD configuration for specific tasks and model scales.